Find the value of a
1/a = 1/3 + 1/4 + 1/6

Answer by edjones(8007)

1/a = 1/3 + 1/4 + 1/6
12=4a+3a+2a multiply each side by 12a
9a=12
a=12/9=4/3
.
Ed